# Thermo-kinetic invariant in anisotropic plasma with heat fluxes

### v1

@@ Abstract
We investigate linear perturbations in an incompressible, weakly collisional, anisotropic plasma, taking heat fluxes into account. Our analysis identifies a new class of time-invariant perturbations that encompass heat flux variations and aperiodic longitudinal perturbations in the magnetic field and flow velocity. Under the Chew-Goldberger-Low (CGL) approximation, which neglects low-frequency heat flux effects, these perturbations are purely thermal and decoupled from magnetic and velocity fields. When heat fluxes are included, these perturbations gain kinematic and magnetic components, leading to the identification of a “thermo-kinetic mode.”
The nature of aperiodic linear modes suggests that they may reach higher amplitudes in perturbation structures, that can be revealed as filamentary formations. We discuss the potential implications of these filaments on the large-scale dynamics of flows and on the nonlinear energy balance within small-scale MHD turbulence. Our findings indicate that thermo-kinetic invariant may play a significant role in modulating turbulence energy transfer dynamics, potentially impacting both local and global flow behavior in weakly collisional anisotropic plasmas.
@@ I Introduction
The paper is organized as follows: In Section II, we present the governing equations for a weakly collisional, anisotropic plasma, incorporating both pressure anisotropy and heat fluxes. We derive the linearized perturbation equations and analyze the conditions under which the thermo-kinetic invariant emerges. Section III summarizes the key findings of the paper and discusses potential applications of these results to the nonlinear global dynamics of plasma flows in various contexts.
@@ II.2 Linear Perturbations
The dynamics of linear perturbations over the above described equilibrium state may be analyzed using the standard linearization procedure. For this we introduce generalized vectors representing equilibrium () and perturbed () components of the physical variables:
@@
Thus, by introducing linear perturbations over the background values and neglecting the nonlinear terms, we can derive a system of linear differential equations that describe the dynamics of perturbations in the anisotropic medium. By utilizing generalized vectors, we can represent the dynamical equations in matrix form as follows:
@@ II.3 Stationary Solution
It appears that the perturbations of the linear stationary mode encompass heat fluxes in both parallel and perpendicular directions, as well as perturbations in velocity and magnetic field along the background magnetic field. Consequently, we can identify as the thermo-kinetic invariant of the linear system. Existence of such conserved quantity highlights the presence of stable, non-oscillatory modes, which may play a significant role in the redistribution of energy within weakly collisional anisotropic plasmas.
@@
Based on the definitions provided above, we can express the thermo-kinetic invariant in the presence of heat fluxes () in a manner that highlights its magnetic and kinetic components as follows:
@@
cannot excite the thermo-kinetic mode with a notable amplitude. As a result, perturbation field of the invariant can be characterized as localized filaments, which lead to the excitation of an aperiodic thermo-kinetic quantity that is conserved throughout the time evolution of the anisotropic flow.
@@ III Summary
In the CGL approximation, this mode is purely thermal and decoupled from magnetic and velocity fields. However, when the heat fluxes are explicitly incorporated into the governing equations, we identify a new class of invariant perturbations with both kinematic and magnetic field components, which we refer to as the “thermo-kinetic mode.” This new linear mode reveals an important link between thermal and kinetic aspects of the plasma, indicating that the heat fluxes, often neglected in simpler models, plays a significant role in shaping the plasma dynamics, particularly in weakly collisional, anisotropic flows.
@@
Our analysis shows that these thermo-kinetic modes persist as aperiodic, invariant quantities that may have profound implications for energy transfer and stability in weakly collisional plasma environments. These modes could influence the long-term behavior of turbulent plasma flows by redistributing energy between different degrees of freedom, including thermal, magnetic, and velocity fields.
@@
We have derived the perturbation field of the thermo-kinetic mode in the coordinate space. It seems that the thermo-kinetic invariant is characterized by small-scale structures, such as filaments, within the perturbation field. This property may significantly alter the energy balance of turbulent flows and resulting turbulent spectrum in weakly collisional plasma flows.
@@
In small-scale MHD turbulence, the governing physical processes exhibit significantly more diversity than those observed at larger scales[19, 20, 21, 22]. The linear conservative modes may add to this complexity: the mode effectively emerging at smaller scales suggests a shift in the nonlinear balance within the turbulence spectrum, indicating that the energy dynamics at these scales may be fundamentally altered. The MHD turbulence cascade is normally dominated by nonlinear couplings between different linear modes. These nonlinear interactions drive energy transfer between modes and a new aperiodic mode can give rise to a new energy exchange channels in the flow. At small scales, perturbations of the linear thermo-kinetic mode can potentially shift how energy is distributed across scales and redefine how turbulent energy dissipate. In this case, thermal effects can unexpectedly produce significant measurable impact, in addition to the classical dissipative effects. Thus, a new aspects of small-scale MHD turbulence dynamics that are essential for understanding energy transport and magnetic field evolution in astrophysical and laboratory plasma environments can be studied.
@@
It remains to be determined whether the thermo-kinetic mode is purely a linear microscopic characteristic of perturbations or if it can also manifest at macroscopic scales. If the latter is true, it could lead to a direct connection between thermal perturbations and the longitudinal kinetic energy of velocity perturbations. This connection could substantially affect the macroscopic acceleration and deceleration of anisotropic magnetized plasma flows, a phenomenon that may have significant implications for the dynamics of solar and stellar wind flows.

### v2

# Thermo-mechanic invariant in anisotropic plasma with heat fluxes

@@ Abstract
We investigate linear perturbations in an incompressible, weakly collisional, anisotropic plasma, taking heat fluxes into account. Our analysis identifies a new class of time-invariant perturbations that encompass heat flux variations and aperiodic longitudinal perturbations in the magnetic field and flow velocity. Under the Chew-Goldberger-Low (CGL) approximation, which neglects low-frequency heat flux effects, these perturbations are purely thermal and decoupled from magnetic and velocity fields. When heat fluxes are included, these perturbations gain kinematic and magnetic components, leading to the identification of a “thermo-mechanic mode.”
The nature of aperiodic linear modes suggests that they may reach higher amplitudes in perturbation structures, that can be revealed as filamentary formations. We discuss the potential implications of these filaments on the large-scale dynamics of flows and on the nonlinear energy balance within small-scale MHD turbulence. Our findings indicate that thermo-mechanic invariant may play a significant role in modulating turbulence energy transfer dynamics, potentially impacting both local and global flow behavior in weakly collisional anisotropic plasmas.
@@ I Introduction
The paper is organized as follows: In Section II, we present the governing equations for a weakly collisional, anisotropic plasma, incorporating both pressure anisotropy and heat fluxes. We derive the linearized perturbation equations and analyze the conditions under which the thermo-mechanic invariant emerges. Section III summarizes the key findings of the paper and discusses potential applications of these results to the nonlinear global dynamics of plasma flows in various contexts.
@@ II.2 Linear Perturbations
The dynamics of linear perturbations over the above described equilibrium state may be analyzed using the standard linearization procedure. For the shortness of the notations we introduce a generalized vector for the physical variables of the problem:
Hence, we may introduce generalized vectors representing equilibrium () and perturbed () components of the physical variables:
Having in mind the equilibrium state introduced by the physical model of our problem we derive:
Thus, by introducing linear perturbations over the background values and neglecting the nonlinear terms, we can derive a system of linear differential equations that describe the dynamics of perturbations in the anisotropic medium. To represent the system in the non-dimensional form we utilize the generalized vector with normalized physical variables
@@
and introduce the typical length-scale of the problem and switch to the non-dimensional description using:
Hence, the dynamical equations can be represented in the non-dimensional matrix form as follows:
@@
The non-stationary solutions for linear perturbation system of incompressible anisotropic flows is already analyzed using the 6 order dispersion equation[5]. Our eigenfunction analysis reveals that the 7th order matrix can describe an additional mode, which does not appear in the time variation analysis but can possess a non-zero, non-trivial solution.
@@ II.3 Stationary Solution
It appears that the perturbations of the linear stationary mode encompass heat fluxes in both parallel and perpendicular directions, as well as perturbations in velocity and magnetic field along the background magnetic field. Consequently, we can identify as the thermo-mechanic invariant of the linear system. Existence of such conserved quantity highlights the presence of stable, non-oscillatory modes, which may play a significant role in the redistribution of energy within weakly collisional anisotropic plasmas.
@@
Based on the definitions provided above, we can express the thermo-mechanic invariant in the presence of heat fluxes () in a manner that highlights its magnetic and kinetic components as follows:
@@
cannot excite the thermo-mechanic mode with a notable amplitude. As a result, perturbation field of the invariant can be characterized as localized filaments, which lead to the excitation of an aperiodic thermo-mechanic quantity that is conserved throughout the time evolution of the anisotropic flow.
To show an illustrative example of the thermo-mechanic invariant field we use 2D reduction to the plane. In this limit heat fluxes contribute into the thermal component of invariant as follows:
Assuming simplest heat flux perturbation field with power low spacial distribution contributing to the uniform in space invariant we may construct solution of the Eq. (48) as follows:
Figure 1 illustrates a specific configuration of the heat flux field of the thermo-mechanic invariant for a given set of parameters. The geometry of heat fluxes in the aperiodic mode appears to be strongly influenced by the anisotropy parameter .
@@ III Summary
In the CGL approximation, this mode is purely thermal and decoupled from magnetic and velocity fields. However, when the heat fluxes are explicitly incorporated into the governing equations, we identify a new class of invariant perturbations with both kinematic and magnetic field components, which we refer to as the “thermo-mechanic mode.” This new linear mode reveals an important link between thermal and kinetic aspects of the plasma, indicating that the heat fluxes, often neglected in simpler models, plays a significant role in shaping the plasma dynamics, particularly in weakly collisional, anisotropic flows.
@@
Our analysis shows that these thermo-mechanic modes persist as aperiodic, invariant quantities that may have profound implications for energy transfer and stability in weakly collisional plasma environments. These modes could influence the long-term behavior of turbulent plasma flows by redistributing energy between different degrees of freedom, including thermal, magnetic, and velocity fields.
@@
We have derived the perturbation field of the thermo-mechanic mode in the coordinate space. It seems that the thermo-mechanic invariant is characterized by small-scale structures, such as filaments, within the perturbation field. This property may significantly alter the energy balance of turbulent flows and resulting turbulent spectrum in weakly collisional plasma flows.
@@
In small-scale MHD turbulence, the governing physical processes exhibit significantly more diversity than those observed at larger scales[19, 20, 21, 22]. The linear conservative modes may add to this complexity: the mode effectively emerging at smaller scales suggests a shift in the nonlinear balance within the turbulence spectrum, indicating that the energy dynamics at these scales may be fundamentally altered. The MHD turbulence cascade is normally dominated by nonlinear couplings between different linear modes. These nonlinear interactions drive energy transfer between modes and a new aperiodic mode can give rise to a new energy exchange channels in the flow. At small scales, perturbations of the linear thermo-mechanic mode can potentially shift how energy is distributed across scales and redefine how turbulent energy dissipate. In this case, thermal effects can unexpectedly produce significant measurable impact, in addition to the classical dissipative effects. Thus, a new aspects of small-scale MHD turbulence dynamics that are essential for understanding energy transport and magnetic field evolution in astrophysical and laboratory plasma environments can be studied.
@@
It remains to be determined whether the thermo-mechanic mode is purely a linear microscopic characteristic of perturbations or if it can also manifest at macroscopic scales. If the latter is true, it could lead to a direct connection between thermal perturbations and the longitudinal kinetic energy of velocity perturbations. This connection could substantially affect the macroscopic acceleration and deceleration of anisotropic magnetized plasma flows, a phenomenon that may have significant implications for the dynamics of solar and stellar wind flows.